Safia bought 3 books and 5 pens for RS60. Each book costs 5 times as much as the pen,
find the cost of 2 books and 7 pens.
1
Expert's answer
2011-06-14T09:46:46-0400
Safia bought 3 books and 5 pens for RS60. Each book costs 5 times as much& as& the& pen,& find the cost of 2 books and 7 pens.
Let 1 book cost a Rs and 1 pen cost b RS. Safia bought 3 books and 5 pens for RS60. So,
3*a+5*b=60.
Each book costs 5 times as much& as the pen, so
a=5*b.
So we get system
a=5*b
3*a+5*b=60

3*(5*b)+5*b=60 -> 20*b=60 -> b=60/20=3 RS.

Then 1 book cost
a=5*b=5*3=15 RS.